OpenStudy (anonymous):

What is the slope of a line perpendicular to the line whose equation is y = − 2/7 x − 5? A. -2/7 B. -7/2 C. 2/7 D. 7/2

OpenStudy (anonymous):

Answer is D; 7/2

OpenStudy (anonymous):

to find the slope of a perpendicular, you first have to find the slope of the equation of the line given. y = -2/7x - 5. In y = mx + b form, the slope is in the m position. So the slope of the line given is -2/7. To find a perpendicular slope, you need to get the negative reciprocal of the slope given. All that means is " flip " the slope and change the sign. So the negative reciprocal of -2/7 is 7/2 (see how I flipped the slope and changed the sign. Karebear is correct, I just thought I might explain it to you so you can understand.